# Enhancing Performance of Explainable AI Models with Constrained Concept Refinement

### v1

@@ 3 Our Proposed Framework
There are two critical challenges to address before answering this question. First, modifying ${\bm{D}}$ may diminish the encoded information that is interpretable by humans, thereby compromising the interpretable-by-design principle of information pursuit. Second, treating the query features as variables introduces significant complexity in deriving the optimal decision rule, as the closed-form expression of IP-OMP presented in Theorem 1 is no longer valid.
@@
To motivate the choice of the surrogate estimator $\tilde{f}_{{\bm{D}}}$, let us revisit the generative model in 1, and additionally assume that the ground truth latent feature matrix ${\bm{D}}^{*}$ is column-orthogonal.
By assuming ${\bm{D}}^{*}$ to be column-orthogonal, we effectively assume an ideal scenario in which queries are mutually independent. Specifically, it follows that $\langle{\mathbf{d}}^{*}_{i},{\mathbf{z}}\rangle$ is independent of $\langle{\mathbf{d}}^{*}_{j},{\mathbf{z}}\rangle$ when ${\mathbf{d}}^{*}_{i}\perp{\mathbf{d}}^{*}_{j}$ and ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{d\times d})$. The primary motivation for focusing on an column-orthogonal ${\bm{D}}^{*}$ is that, under this assumption, the estimator $f_{{\bm{D}}}$ obtained from IP-OMP becomes inherently differentiable with respect to ${\bm{D}}$:
Lemma 1.
@@
The proof of Lemma 1 is deferred to Section C.4.1. Assuming a small initial error, i.e., $\|{\bm{D}}^{*}-{\bm{D}}\|\leq\rho$ for some small $\rho>0$, the iterates ${\bm{D}}^{(t)}={\bm{D}}+\Delta{\bm{D}}^{(t)}$ of Algorithm 1 stay approximately column-orthogonal. This consideration motivates the introduction of the following surrogate estimator $\tilde{f}$:
@@
It is important to note that we do not require $\tilde{f}$ to be differentiable everywhere; rather, differentiability is required only almost surely along the trajectory of Algorithm 1, which we will rigorously demonstrate. Alternative differentiable surrogates for $f$ include the task-driven dictionary learning method [36] and the unrolled dictionary learning method [37, 57], which can be used in place of Section 3.
@@
$\Delta{\bm{D}}^{(t+1)}=\operatorname*{arg\,min}_{\|\Delta{\bm{D}}\|_{1,2}\leq\rho}\!\left\|\Delta{\bm{D}}\!-\!\left(\Delta{\bm{D}}^{(t)}\!-\!\eta\frac{\partial\mathcal{L}(\tilde{f}_{{\bm{D}}+\Delta{\bm{D}}^{(t)}})}{\partial\Delta{\bm{D}}^{(t)}}\right)\right\|_{2},$
@@ 3.1 Accuracy and Interpretability Guarantees
Assumption 3 (Probabilistic generative model).
Suppose that the target random variable $y$ and queries are generated according to Assumption 1. Moreover, suppose that ground truth latent feature matrix ${\bm{D}}^{*}$ is column-orthogonal and the input ${\mathbf{x}}$ satisfies ${\mathbf{x}}={\bm{D}}^{*}\mathbf{\beta}$, where: (1) the support of $\mathbf{\beta}$, denoted as $S^{*}$, is drawn uniformly from the set of $k$-element subsets of $[n]$; and (2) the non-zero elements of $\mathbf{\beta}$ are i.i.d. and satisfy $\mathbb{E}[\mathbf{\beta}_{i}]=0$, $\mathrm{Var}[\mathbf{\beta}_{i}]=\sigma^{2}$, and $\gamma\leq|\mathbf{\beta}_{i}|\leq\Gamma$.
The generative model in 3 is akin to those explored in the sparse coding literature [5, 2, 45, 34]. Under this model, prediction error can be evaluated using the squared loss, while interpretability can be assessed by the proximity of the query feature matrix ${\bm{D}}^{(T)}$, produced by CCR, to the ground truth latent feature matrix ${\bm{D}}^{*}$.
Our next theorem shows that CCR effectively resolves the challenge faced by IP-OMP, as outlined in Theorem 2.111We note that the result of Theorem 2 also holds for the generative model described in 3. Let ${\bm{D}}^{(t)}={\bm{D}}+\Delta{\bm{D}}^{(t)}$ be the corrected query feature matrix generated by CCR at iteration $t$.
Theorem 3.
Suppose that 3 holds. Moreover, suppose that that the initial query feature matrix ${\bm{D}}$ satisfies $\|{\bm{D}}-{\bm{D}}^{*}\|_{1,2}=\rho\leq\frac{\gamma}{8\sqrt{k}\Gamma}$ and the step-size $\eta$ satisfies $0<\eta<\frac{1}{2\|{\mathbf{x}}\|^{2}_{2}}$. Algorithm 1 with surrogate estimator (3) and update rule (3) satisfies:
@@
While Theorem 3 addresses the performance limitations of IP-OMP, it does not guarantee improved interpretability. This outcome is unsurprising, since the reliance of the method on only a single input ${\mathbf{x}}$ ensures that only the columns of ${\bm{D}}$ with indices in $S^{*}$ can be improved, essentially leaving the columns outside $S^{*}$ unmodified. Moreover, if ${\bm{D}}^{*}$ is not full-rank, our analysis shows that ${\bm{D}}$ may deviate from ${\bm{D}}^{*}$ in directions column-orthogonal to the column space of ${\bm{D}}^{*}$; these deviations cannot be eliminated by Section 3. This observation is further supported by numerical experiments (see Figure 4).
To tackle these challenges, we assume that ${\bm{D}}^{*}$ is full-rank. Moreover, it is essential to use a sufficient number of i.i.d. input samples to ensure that each column of ${\bm{D}}^{*}$ contributes to at least one input sample. This aligns more closely with practical scenarios, such as image classification, where multiple samples are typically used during the training phase. We denote the i.i.d. input samples and their corresponding target random variables, generated according to 3, as $\{{\mathbf{x}}^{h}\}_{h=1}^{m}$ and $\{y^{h}\}_{h=1}^{m}$, respectively. For any estimator ${\mathbf{y}}_{\mathrm{pred}}=\{y_{\mathrm{pred}}^{h}\}_{h=1}^{m}$, the aggregated squared loss is defined as
@@
Theorem 4.
Suppose that $\{{\mathbf{x}}^{h}\}^{m}_{h=1}$ and $\{y^{h}\}^{m}_{h=1}$ are i.i.d. samples generated from 3 with a fixed full-rank column-orthogonal ${\bm{D}}^{*}$. Suppose that $m=\Omega\left(\frac{n^{6}}{\sigma^{2}k^{5}}\right)$. Moreover, suppose that the initial query feature matrix ${\bm{D}}$ satisfies $\|{\bm{D}}-{\bm{D}}^{*}\|_{1,2}=\rho\leq\frac{\gamma}{8\sqrt{k}\Gamma}$ and the step-size $\eta$ satisfies $\eta=O\left(\frac{1}{\sigma^{2}}\right)$. With high probability, Algorithm 1 with surrogate estimator (3) and update rule (3) applied to the aggregated squared loss (5) satisfies:
@@
According to Theorem 4, by leveraging multiple samples, CCR achieves the best of both worlds: it guarantees convergence to the ground truth latent feature matrix ${\bm{D}}^{*}$, progressively enhancing interpretability, while simultaneously driving $\mathcal{L}_{m}$ to zero, thereby achieving perfect accuracy.
Contrary to traditional analyses in dictionary learning [5, 34], the proofs of Theorem 3 and Theorem 4 require a careful alignment of the updates from gradient descent with the direction of the true solution. In the setting of Theorem 4, this is even more challenging because the finite sample size $m$ inevitably causes deviations from the population-level behavior. Moreover, we must further control the deviation introduced by the projection step. The detailed proofs of these theorems are provided in Appendix C. We also include a detailed discussion on both theorems in Appendix B and experiments that validate them in Section D.1.
@@ 4 Application: Interpretable Image Classification
The image classification setting differs slightly from the setting provided in Definition 1. Specifically, we are given a dataset $\{{\mathbf{a}}^{i},{\mathbf{y}}^{i}\}_{i=1}^{m}$ consisting of $m$ image (${\mathbf{a}}$)-label (${\mathbf{y}}$) pairs, where the goal is to predict the correct label for a given image. Here, ${\mathbf{y}}$ represents a one-hot label vector. Additionally, this setting provides access to a concept set $\{{\mathbf{c}}_{i}\}_{i=1}^{n}$, which consists of $n$ key concepts that can aid in classifying an image. These concepts are typically textual; for example, in the case of animal images, the concept set might include descriptors like “stocky body” or “black stripes”.
@@
CLIP embedding. We employ the multi-modal model CLIP [44]—a recently introduced large Vision-Language Model—to embed both images and concepts into the same latent space.
@@
Embedding normalization and projection. After each iteration, we perform a projection step similar to Section 3 for each updated concept embedding. Additionally, we normalize each embedding to ensure that it remains on the unit ball (see Algorithm 4 in Section D.2).
@@ 4.1 Performance
We compare the performance of Algorithm 2 against two recently proposed explainable AI methods, CLIP-IP-OMP [14] and label-free CBM (lf-CBM) [40], as well as the CCR baseline without the concept refinement step, in the context of explainable image classification.
@@
The evaluation is conducted across five image classification benchmarks: CIFAR-10, CIFAR-100 [26], ImageNet [16], CUB-200 [62], and Places365 [68].
Figure 2: Prediction accuracy of CCR and its baseline across iterations, with the final test accuracy of CLIP-IP-OMP and lf-CBM indicated for reference.
To ensure a fair comparison, all methods use the same concept set, which is generated by GPT-3 [9]. A detailed description of the generation process can be found in [40]. For the CIFAR-10/100 and CUB-200 datasets, we tune CLIP-IP-OMP to match the average sparsity level of ${\mathbf{s}}$, also referred to as the explanation length or $k$, used in CCR. For ImageNet and Places365, we report the best accuracy achieved by CLIP-IP-OMP across all explanation lengths. Since Lf-CBM does not allow tuning of its explanation length, we directly report its accuracy. The constraint parameter $\rho$ for CCR is fixed at $0.1$ for all experiments.
@@
Thanks to the parallelization enabled by hard thresholding and the computational efficiency of backpropagation, CCR significantly reduces computational costs compared to CLIP-IP-OMP and lf-CBM, particularly when applied to large-scale datasets such as ImageNet ($\approx 1.2$ million images) and Places365 ($\approx 1.8$ million images). As reported in [14], training lf-CBM on ImageNet requires $\approx 50$ hours on an NVIDIA RTX A5000 GPU. Under the same experimental conditions, CLIP-IP-OMP, with an average explanation length of $\approx 50$, incurs a computational cost of $\approx 40$ hours. In our computational environment, using an NVIDIA Tesla V100 GPU, CLIP-IP-OMP remains comparably expensive, requiring $\approx 33$ hours for $k=50$. In contrast, CCR (with $k\approx 49$, as shown in Table 1) processes ImageNet in only $\approx 2$ hours (corresponding to 200 iterations) while achieving even higher accuracy, demonstrating a substantial improvement in computational efficiency. To make CLIP-IP-OMP more computationally feasible, one can reduce $k$; however, for $k=10$, while the processing time drops to $\approx 6$ hours, it comes at the cost of a significant accuracy decline to $\approx 63\%$.
@@ 4.2 Interpretability
To illustrate the interpretability of Algorithm 2, we follow the methodology outlined in [14] and present the most significant coefficients and weights from the algorithm on different samples from the Places365 dataset. As shown in Figure 3, we highlight the top 10 concepts in ${\mathbf{s}}$ with the highest values (left) along with their corresponding weights in the linear layer ${\mathbf{L}}$ for the predicted label (right). The selected concepts are semantically relevant to the image, and the linear layer effectively assigns substantial weight to key concepts such as “designated fairways and greens”, enabling accurate prediction while appropriately disregarding concepts that, although relevant, may be misleading, such as “a flag” or “a country”. For additional case studies on other datasets, we refer the reader to Section D.3.

### v2

@@ 3 Our Proposed Framework
There are two critical challenges to address before answering this question. First, modifying ${\bm{D}}$ may diminish the encoded information that is interpretable by humans, thereby compromising the interpretable-by-design principle of information pursuit. Second, treating the query features as variables introduces significant complexity in deriving the optimal decision rule, as the closed-form expression of IP-OMP presented in Theorem 2.5 is no longer valid.
@@
To motivate the choice of the surrogate estimator $\tilde{f}_{{\bm{D}}}$, let us revisit the generative model in 2.3, and additionally assume that the ground truth latent feature matrix ${\bm{D}}^{*}$ is column-orthogonal. By assuming ${\bm{D}}^{*}$ to be column-orthogonal, we effectively assume an ideal scenario in which queries are mutually independent. Specifically, it follows that $\langle{\mathbf{d}}^{*}_{i},{\mathbf{z}}\rangle$ is independent of $\langle{\mathbf{d}}^{*}_{j},{\mathbf{z}}\rangle$ when ${\mathbf{d}}^{*}_{i}\perp{\mathbf{d}}^{*}_{j}$ and ${\mathbf{z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{d\times d})$. The primary motivation for focusing on an column-orthogonal ${\bm{D}}^{*}$ is that, under this assumption, the estimator $f_{{\bm{D}}}$ obtained from IP-OMP becomes inherently differentiable with respect to ${\bm{D}}$:
Lemma 3.1.
@@
The proof of Lemma 3.1 is deferred to Section C.4.1. Assuming a small initial error, i.e., $\|{\bm{D}}^{*}-{\bm{D}}\|\leq\rho$ for some small $\rho>0$, the iterates ${\bm{D}}^{(t)}={\bm{D}}+\Delta{\bm{D}}^{(t)}$ of Algorithm 1 stay approximately column-orthogonal. This consideration motivates the introduction of the following surrogate estimator $\tilde{f}$:
@@
It is important to note that we do not require $\tilde{f}$ to be differentiable everywhere; rather, differentiability is required only almost surely along the trajectory of Algorithm 1, which we will rigorously demonstrate. Alternative differentiable surrogates for $f$ include the task-driven dictionary learning method (Mairal et al., 2011) and the unrolled dictionary learning method (Malézieux et al., 2021; Tolooshams & Ba, 2021), which can be used in place of Section 3.
@@
$\Delta{\bm{D}}^{(t+1)}\\ =\!\operatorname*{arg\,min}_{\|\Delta{\bm{D}}\|_{1,2}\leq\rho}\!\left\|\Delta{\bm{D}}\!-\!\left(\Delta{\bm{D}}^{(t)}\!-\!\eta\frac{\partial\mathcal{L}(\tilde{f}_{{\bm{D}}+\Delta{\bm{D}}^{(t)}})}{\partial\Delta{\bm{D}}^{(t)}}\right)\right\|_{2},$ (5)
@@ 3.1 Accuracy and Interpretability Guarantees
Assumption 3.2 (Probabilistic generative model).
Suppose that the target random variable $y$ and queries are generated according to Assumption 2.3. Moreover, suppose that ground truth latent feature matrix ${\bm{D}}^{*}$ is column-orthogonal and the input ${\mathbf{x}}$ satisfies ${\mathbf{x}}={\bm{D}}^{*}\bm{\beta}$, where: (1) the support of $\bm{\beta}$, denoted as $S^{*}$, is drawn uniformly from the set of $k$-element subsets of $[n]$; and (2) the non-zero elements of $\bm{\beta}$ are i.i.d. and satisfy $\mathbb{E}[\bm{\beta}_{i}]=0$, $\mathrm{Var}[\bm{\beta}_{i}]=\sigma^{2}$, and $\gamma\leq|\bm{\beta}_{i}|\leq\Gamma$.
The generative model in 3.2 is akin to those explored in the sparse coding literature (Arora et al., 2015; Agarwal et al., 2016; Ravishankar et al., 2020; Liang et al., 2022). Under this model, prediction error can be evaluated using the squared loss, while interpretability can be assessed by the proximity of the query feature matrix ${\bm{D}}^{(T)}$, produced by CCR, to the ground truth latent feature matrix ${\bm{D}}^{*}$.
Our next theorem shows that CCR effectively resolves the challenge faced by IP-OMP, as outlined in Theorem 2.6.111We note that the result of Theorem 2.6 also holds for the generative model described in 3.2. Let ${\bm{D}}^{(t)}={\bm{D}}+\Delta{\bm{D}}^{(t)}$ be the corrected query feature matrix generated by CCR at iteration $t$.
Theorem 3.3.
Suppose that 3.2 holds. Moreover, suppose that that the initial query feature matrix ${\bm{D}}$ satisfies $\|{\bm{D}}-{\bm{D}}^{*}\|_{1,2}=\rho\leq\frac{\gamma}{8\sqrt{k}\Gamma}$ and the step-size $\eta$ satisfies $0<\eta<\frac{1}{2\|{\mathbf{x}}\|^{2}_{2}}$. Algorithm 1 with surrogate estimator (3) and update rule (5) satisfies:
@@
While Theorem 3.3 addresses the performance limitations of IP-OMP, it does not guarantee improved interpretability. This outcome is unsurprising, since the reliance of the method on only a single input ${\mathbf{x}}$ ensures that only the columns of ${\bm{D}}$ with indices in $S^{*}$ can be improved, essentially leaving the columns outside $S^{*}$ unmodified. Moreover, if ${\bm{D}}^{*}$ is not full-rank, our analysis shows that ${\bm{D}}$ may deviate from ${\bm{D}}^{*}$ in directions column-orthogonal to the column space of ${\bm{D}}^{*}$; these deviations cannot be eliminated by Equation 5. This observation is further supported by numerical experiments (see Figure 4).
To tackle these challenges, we assume that ${\bm{D}}^{*}$ is full-rank. Moreover, it is essential to use a sufficient number of i.i.d. input samples to ensure that each column of ${\bm{D}}^{*}$ contributes to at least one input sample. This aligns more closely with practical scenarios, such as image classification, where multiple samples are typically used during the training phase. We denote the i.i.d. input samples and their corresponding target random variables, generated according to 3.2, as $\{{\mathbf{x}}^{h}\}_{h=1}^{m}$ and $\{y^{h}\}_{h=1}^{m}$, respectively. For any estimator ${\mathbf{y}}_{\mathrm{pred}}=\{y_{\mathrm{pred}}^{h}\}_{h=1}^{m}$, the aggregated squared loss is defined as
@@
Theorem 3.4.
Suppose that $\{{\mathbf{x}}^{h}\}^{m}_{h=1}$ and $\{y^{h}\}^{m}_{h=1}$ are i.i.d. samples generated from 3.2 with a fixed full-rank column-orthogonal ${\bm{D}}^{*}$. Suppose that $m=\Omega\left(\frac{n^{6}}{\sigma^{2}k^{5}}\right)$. Moreover, suppose that the initial query feature matrix ${\bm{D}}$ satisfies $\|{\bm{D}}-{\bm{D}}^{*}\|_{1,2}=\rho\leq\frac{\gamma}{8\sqrt{k}\Gamma}$ and the step-size $\eta$ satisfies $\eta=O\left(\frac{1}{\sigma^{2}}\right)$. With high probability, Algorithm 1 with surrogate estimator (3) and update rule (5) applied to the aggregated squared loss (6) satisfies:
@@
According to Theorem 3.4, by leveraging multiple samples, CCR achieves the best of both worlds: it guarantees convergence to the ground truth latent feature matrix ${\bm{D}}^{*}$, progressively enhancing interpretability, while simultaneously driving $\mathcal{L}_{m}$ to zero, thereby achieving perfect accuracy.
Contrary to traditional analyses in dictionary learning (Arora et al., 2015; Liang et al., 2022), the proofs of Theorem 3.3 and Theorem 3.4 require a careful alignment of the updates from gradient descent with the direction of the true solution. In the setting of Theorem 3.4, this is even more challenging because the finite sample size $m$ inevitably causes deviations from the population-level behavior. Moreover, we must further control the deviation introduced by the projection step. The detailed proofs of these theorems are provided in Appendix C. We also include a detailed discussion on both theorems in Appendix B and experiments that validate them in Section D.1.
@@ 4 Application: Interpretable Image Classification
The image classification setting differs slightly from the setting provided in Definition 2.1. Specifically, we are given a dataset $\{{\mathbf{a}}^{i},{\mathbf{y}}^{i}\}_{i=1}^{m}$ consisting of $m$ image (${\mathbf{a}}$)-label (${\mathbf{y}}$) pairs, where the goal is to predict the correct label for a given image. Here, ${\mathbf{y}}$ represents a one-hot label vector. Additionally, this setting provides access to a concept set $\{{\mathbf{c}}_{i}\}_{i=1}^{n}$, which consists of $n$ key concepts that can aid in classifying an image. These concepts are typically textual; for example, in the case of animal images, the concept set might include descriptors like “stocky body” or “black stripes”.
@@
CLIP embedding. We employ the multi-modal model CLIP (Radford et al., 2021)—a recently introduced large Vision-Language Model—to embed both images and concepts into the same latent space.
@@
Embedding normalization and projection. After each iteration, we perform a projection step similar to Equation 5 for each updated concept embedding. Additionally, we normalize each embedding to ensure that it remains on the unit ball (see Algorithm 4 in Section D.2).
@@ 4.1 Performance
We compare the performance of Algorithm 2 against two recently proposed explainable AI methods, CLIP-IP-OMP (Chattopadhyay et al., 2024) and label-free CBM (lf-CBM) (Oikarinen et al., 2023), as well as the CCR baseline without the concept refinement step, in the context of explainable image classification.
@@
The evaluation is conducted across five image classification benchmarks: CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), ImageNet (Deng et al., 2009), CUB-200 (Wah et al., 2011), and Places365 (Zhou et al., 2017).
Figure 2: Prediction accuracy of CCR and its baseline across iterations, with the final test accuracy of CLIP-IP-OMP and lf-CBM indicated for reference. For CCR and its baseline, we run each experiment for five times and present the average test accuracy at each time step. The shaded area is bounded by the maximum and minimum accuracy obtained over five runs.
To ensure a fair comparison, all methods use the same concept set, which is generated by GPT-3 (Brown et al., 2020). A detailed description of the generation process can be found in (Oikarinen et al., 2023). For the CIFAR-10/100 and CUB-200 datasets, we tune CLIP-IP-OMP to match the average sparsity level of ${\mathbf{s}}$, also referred to as the explanation length or $k$, used in CCR. For ImageNet and Places365, we report the best accuracy achieved by CLIP-IP-OMP across all explanation lengths. Since Lf-CBM does not allow tuning of its explanation length, we directly report its accuracy. The constraint parameter $\rho$ for CCR is fixed at $0.1$ for all experiments.
@@
Thanks to the parallelization enabled by hard thresholding and the computational efficiency of backpropagation, CCR significantly reduces computational costs compared to CLIP-IP-OMP and lf-CBM, particularly when applied to large-scale datasets such as ImageNet ($\approx 1.2$ million images) and Places365 ($\approx 1.8$ million images). As reported in (Chattopadhyay et al., 2024), training lf-CBM on ImageNet requires $\approx 50$ hours on an NVIDIA RTX A5000 GPU. Under the same experimental conditions, CLIP-IP-OMP, with an average explanation length of $\approx 50$, incurs a computational cost of $\approx 40$ hours. In our computational environment, using an NVIDIA Tesla V100 GPU, CLIP-IP-OMP remains comparably expensive, requiring $\approx 33$ hours for $k=50$. In contrast, CCR (with $k\approx 49$, as shown in Table 1) processes ImageNet in only $\approx 2$ hours (corresponding to 200 iterations) while achieving even higher accuracy, demonstrating a substantial improvement in computational efficiency. To make CLIP-IP-OMP more computationally feasible, one can reduce $k$; however, for $k=10$, while the processing time drops to $\approx 6$ hours, it comes at the cost of a significant accuracy decline to $\approx 63\%$.
@@ 4.2 Interpretability
To illustrate the interpretability of Algorithm 2, we follow the methodology outlined in (Chattopadhyay et al., 2024) and present the most significant coefficients and weights from the algorithm on different samples from the Places365 dataset. As shown in Figure 3, we highlight the top 10 concepts in ${\mathbf{s}}$ with the highest values (left) along with their corresponding weights in the linear layer ${\mathbf{L}}$ for the predicted label (right). The selected concepts are semantically relevant to the image, and the linear layer effectively assigns substantial weight to key concepts such as “designated fairways and greens”, enabling accurate prediction while appropriately disregarding concepts that, although relevant, may be misleading, such as “a flag” or “a country”. For additional case studies on other datasets, we refer the reader to Section D.5.
Impact Statement
This paper introduces Constrained Concept Refinement (CCR), a principled framework that helps bridge the long-standing gap between interpretability and accuracy in machine learning. By constraining the refinement of concept embeddings to lie within a small neighborhood around their initial values, CCR enables interpretable-by-design models to improve prediction performance without compromising their explainability. Our theoretical and empirical results demonstrate the effectiveness of CCR across a variety of tasks and datasets, both in terms of predictive performance and computational efficiency.
The broader impact of this work lies in its potential to advance the practical adoption of interpretable machine learning methods in real-world settings. In particular, the computational efficiency afforded by CCR may facilitate the deployment of explainable artificial intelligence (XAI) techniques in resource-constrained environments.
From an ethical perspective, the capacity to generate explanations that are faithful, stable, and aligned with human-interpretable concepts contributes to addressing critical concerns related to algorithmic accountability and bias. For the proposed method, hyperparameter tuning remains a crucial yet underexplored component for ensuring effective performance in practical applications. This aspect warrants careful consideration to ensure that the resulting outputs are both reliable and justifiable.